OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 5?

OpenStudy (anonymous):

(-3, 4) (1, -2) \[(\sqrt{5},\sqrt{5})\]

ganeshie8 (ganeshie8):

equation of circle centered at origin with radius 5 : \[\large x^2 + y^2 = 5^2\]

ganeshie8 (ganeshie8):

plugin each option into the equation and see which one satisfies

OpenStudy (anonymous):

idk how lol

OpenStudy (anonymous):

@iGreen ill give medal

OpenStudy (igreen):

Graph of a circle centered at origin with radius 5:

OpenStudy (igreen):

Now do you know what the answer is?

OpenStudy (anonymous):

\[\sqrt{5},\sqrt{5}\]?

OpenStudy (igreen):

You can't really see that on the graph. But check -3, 4 and 1, -2 first.

OpenStudy (igreen):

(-3, 4) is on the circle's line..